OpenStudy (katherinkoon):

will medal. Solve the problem and explain how estimation could be used to make sure the answer is reasonable. Olivia bought some bananas for $3.55 and some ice cream for $3.25. About how much change did Olivia get from $10.00?

OpenStudy (katherinkoon):

A. Answer: Olivia got $3.20 in change. Estimation: 3.55 + 3.25 is about 7. $10 – $7 = $3 Olivia got about $3 in change. B. Answer: Olivia got $3.50 in change. Estimation: 3.55 + 3.25 is about 7. $7 ÷ 2 = $3.50 Olivia got about $3.50 in change. C. Answer: Olivia got $16.80 in change. Estimation: 3.55 + 3.25 is about 7. $10 + $7 = $17 Olivia got about $17 in change.

OpenStudy (photon336):

well, I guess you could round up 3.60+3.30 = 6.90~ then subtract this from 10

OpenStudy (anonymous):

The answer is A.